If 23.00 grams of a 45.00 gram sample of unknown isotope remain after 17 days 12 hours, what is the half-life of the unknown isotope to the nearest minute?
Answer by AnlytcPhil(1278)

If 23.00 grams of a 45.00 gram sample of unknown isotope remain after 
17 days 12 hours, what is the half-life of the unknown isotope to the
nearest minute?
The formula is
A = Pert
When time t=0 days, the amount A = 45g. Substitute these two values
45 = Per(0)
45 = Pe0
45 = P(1)
P = 45 grams
so substitute 45 for P in A = Pert
A = 45ert
When time t=17.5 days, the amount A = 23g. Substitute these two values:
23 = 45er(17.5)
23 = 45e17.5r
Divide both sides by 45
23/45 = e17.5r
.5111111111 = e17.5r
Rewrite this equation using the fact that
" N = eM can be rewritten as M = ln(N) "
17.5r = ln(.5111111111)
Use a calculator to find ln(.5111111111) = -.6711682738
17.5r = -.6711682738
Solve for r
r = -.6711682738/17.5
r = -.0383524728
So substitute this value of r in A = 45ert
A = 45e-.0383524728t
We wish to know how many minutes is required for it to become half
its original amount, or 1/2 of 45g, or 22.5 g. So substitute this
for A in A = 45e-.0394804867t
22.5 = 45e-.0383524728t
Divide both sides by 45
22.5/45 = e-.0383524728t
.5 = e-.0383524728t
Rewrite this as
ln(.5) = -.0383524728t
Solve for t
-.6931471806 = -.0383524728t
-.0383524728t = -.6931471806
t = -.6931471806/(-.0383524728)
t = 18.07307665 days
To find hours, multiply .07307665 by 24
18 days 1.753839658 hours
To find minutes, multiply .753839658 by 60
18 days 1 hour 45.23037947 minutes or to nearest minute:
18 days 1 hour, 45 minutes.
Edwin